the length and breadth of a rectangular field is x and y respectively. if x+y=40, then the area of rectangular field is?

Grade:12th pass

1 Answers

Athira maria antony
6 years ago
We know x+y is 40. From this we can find the perimeter== 2(x+y)=2×40 =80And hence u get the perimeter as 80